# A $2$-distance $(2\Delta+7)$-coloring of planar graphs

### v2

@@ 2.3. The case $\Delta=8$
(5.3). Let $m_{3}(v)=6$. Notice first that if $v$ has at most four neighbours forming $4$- or $5$-vertices, then we have $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-4\times\frac{2}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and at most $\frac{2}{9}$ to each of its $5^{-}$-neighbours by R4-R6 and R8-R9. Therefore we may further assume that $v$ has at least five neighbours forming $4$- or $5$-vertices, which clearly infers that $v$ has at least one support vertex. Thus we have $d_{2}(v)\geq 23$ by Lemma 2.38. On the other hand, by Lemma 2.29, we conclude that a $4(1^{+})$-vertex has at most one $4$-neighbour. In addition, a 5(5)555(5)5 ( 5 )-vertex has at most one neighbour consisting of $4$- or 5(5)555(5)5 ( 5 )-vertices by Lemma 2.34(b). This implies that $v$ has at most four 5(5)555(5)5 ( 5 )-neighbours, and at most five $4(1^{+})$-vertices.
@@
Suppose first that $v$ has no 5(5)555(5)5 ( 5 )-neighbours. If $v$ has at most four $4(1^{+})$-neighbours, then $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-4\times\frac{1}{6}-3\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{6}$ to each of its $4(1^{+})$-neighbours by R4-R6, and $\frac{1}{9}$ to each of its $5(4)$-neighbours by R8. If $v$ has five $4(1^{+})$-neighbours, then $v$ cannot have any $5(4)$-neighbours as $d_{2}(v)\geq 23$. So, $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-5\times\frac{1}{6}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{1}{6}$ to each of its $4(1^{+})$-neighbours by R4-R6.
@@
Next we suppose that $v$ has exactly one 5(5)555(5)5 ( 5 )-neighbour. Since $d_{2}(v)\geq 23$, we can say that $v$ has at most four $4$-neighbours. If $v$ has at most two $4$-neighbours, then $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-\frac{2}{9}-2\times\frac{1}{6}-4\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{2}{9}$ to its 5(5)555(5)5 ( 5 )-neighbour by R9(b), $\frac{1}{6}$ to each of its $4(1^{+})$-neighbours by R4-R6, and $\frac{1}{9}$ to each of its $5(4)$-neighbours by R8. If $v$ has at least three $4$-neighbours, then $v$ is adjacent to at most one $5(4)$-vertices as $d_{2}(v)\geq 23$, and so $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-\frac{2}{9}-4\times\frac{1}{6}-\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{2}{9}$ to each of its 5(5)555(5)5 ( 5 )-neighbours by R9(b), $\frac{1}{6}$ to each of its $4(1^{+})$-neighbours by R4-R6, and $\frac{1}{9}$ to its $5(4)$-neighbour by R8.
@@
Suppose now that $v$ has exactly two 5(5)555(5)5 ( 5 )-neighbours. Then we deduce that $v$ has at most three neighbours consisting of $4$- or $5(4)$-vertices as $d_{2}(v)\geq 23$, and so $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-2\times\frac{2}{9}-3\times\frac{1}{6}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{2}{9}$ to each of its 5(5)555(5)5 ( 5 )-neighbours by R9(b), at most $\frac{1}{6}$ to each of its $4(1^{+})$ and $5(4)$-neighbours by R4-R6 and R8.
@@
Suppose that $v$ has exactly three 5(5)555(5)5 ( 5 )-neighbours. It follows that $v$ has at most one $4$-neighbour as $d_{2}(v)\geq 23$ together with Lemma 2.34(a). By the same reason, $v$ has at most two neighbours consisting of $4$- or $5(4)$-vertices. Thus, we have $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-3\times\frac{2}{9}-\frac{1}{6}-\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{2}{9}$ to each of its 5(5)555(5)5 ( 5 )-neighbours by R9(b), $\frac{1}{6}$ to each of its $4(1^{+})$-neighbours by R4-R6, and $\frac{1}{9}$ to each of its $5(4)$-neighbours by R8.
@@
Finally, suppose that $v$ has four 5(5)555(5)5 ( 5 )-neighbours. Then $v$ has no $4$- and $5(4)$-neighbours by Lemma 2.34(a),(b), and so $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-4\times\frac{2}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{2}{9}$ to each of its 5(5)555(5)5 ( 5 )-neighbours by R9(b).
@@
(5.4). Let $m_{3}(v)=7$. By Lemma 2.38, we deduce that $v$ has at most five $5(4^{+})$-neighbours, and so $\mu^{*}(v)\geq 3-7\times\frac{1}{3}-5\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-face by R1, $\frac{1}{9}$ to each of its $5(4^{+})$-neighbour by R8-R9.
@@
(6). Let $k=8$. The initial charge of $v$ is $\mu(v)=d(v)-4=4$. Notice first that if $m_{3}(v)\leq 4$, then we have $\mu^{*}(v)\geq 4-4\times\frac{1}{3}-8\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and at most $\frac{1}{3}$ to each of its neighbours by R3-R10. Therefore, we may assume that $5\leq m_{3}(v)\leq 8$.
@@
(6.1). Let $m_{3}(v)=5$. By Lemma 2.29(c), a $4(3^{+})$-vertex cannot be adjacent to any $4$-vertex. It follows that $v$ has at most five $4(3^{+})$-neighbours. If $v$ has at most four $4(3^{+})$-neighbours, then $\mu^{*}(v)\geq 4-5\times\frac{1}{3}-4\times\frac{1}{3}-4\times\frac{1}{4}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R6(b), R7, and at most $\frac{1}{4}$ to each of its other neighbours by R3-R5 and R8-R10. If $v$ has five $4(3^{+})$-neighbours, then $v$ cannot have any $3$-, $4(2^{-})$- or $5(4^{+})$-neighbour by Lemmas 2.28(a) and 2.29(c). Thus $\mu^{*}(v)\geq 4-5\times\frac{1}{3}-5\times\frac{1}{3}-3\times\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R6(b) and R7, $\frac{1}{18}$ to each of its $6(5)$-neighbours by R10.
(6.2). Let $m_{3}(v)=6$. By Lemma 2.29(c), $v$ has at most five $4(3^{+})$-neighbours. If $v$ has five such neighbours, then $v$ is not adjacent to any $3$-, $4(2^{-})$- or $5(4^{+})$-vertex by Lemmas 2.28(a) and 2.29(c). Thus $\mu^{*}(v)\geq 4-6\times\frac{1}{3}-5\times\frac{1}{3}-3\times\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R6(b) and R7, $\frac{1}{18}$ to each of its $6(5)$-neighbours by R10. We may further assume that $v$ has at most four $4(3^{+})$-neighbours.
@@
Suppose first that $v$ has no $4(3^{+})$-neighbour. Then $\mu^{*}(v)\geq 4-6\times\frac{1}{3}-8\times\frac{1}{4}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and at most $\frac{1}{4}$ to each of its neighbours by R3-R5 and R8-R10.
@@
Next suppose that $v$ has exactly one $4(3^{+})$-neighbour. It follows from Lemmas 2.28(a) and 2.29(c) that $v$ has at most six neighbours consisting of $3$-, $4(2^{-})$- or $5(4^{+})$-vertices. Thus $\mu^{*}(v)\geq 4-6\times\frac{1}{3}-\frac{1}{3}-6\times\frac{1}{4}-\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R6(b) and R7, at most $\frac{1}{4}$ to each of its $3$-, $4(1)$-, $4(2)$- and $5(4^{+})$-neighbours by R3-R5 and R8-R9, $\frac{1}{18}$ to each of its $6(5)$-neighbours by R10.
@@
Suppose now that $v$ has exactly two $4(3^{+})$-neighbours. In such a case, $v$ has at most five neighbours consisting of $3$-, $4(2^{-})$- or $5(4^{+})$-vertices by Lemmas 2.28(a) and 2.29(c). Thus $\mu^{*}(v)\geq 4-6\times\frac{1}{3}-2\times\frac{1}{3}-5\times\frac{1}{4}-\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R6(b) and R7, at most $\frac{1}{4}$ to each of its $3$-, $4(1)$-, $4(2)$- and $5(4^{+})$-neighbours by R3-R5 and R8-R9, $\frac{1}{18}$ to each of its $6(5)$-neighbours by R10.
@@
Suppose that $v$ has exactly three $4(3^{+})$-neighbours. In such a case, $v$ can have at most three neighbours consisting of $3$-, $4(2^{-})$- or $5(4^{+})$-vertices by Lemmas 2.28(a) and 2.29(c). Thus $\mu^{*}(v)\geq 4-6\times\frac{1}{3}-3\times\frac{1}{3}-3\times\frac{1}{4}-2\times\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R6(b) and R7, at most $\frac{1}{4}$ to each of its $3$-, $4(1)$-, $4(2)$- and $5(4^{+})$-neighbours by R3-R5 and R8-R9, $\frac{1}{18}$ to each of its $6(5)$-neighbours by R10.
@@
Finally suppose that $v$ has exactly four $4(3^{+})$-neighbours. In such a case, $v$ can have at most one neighbour consisting of $3$-, $4(2^{-})$- or $5(4^{+})$-vertices by Lemmas 2.28(a) and 2.29(c). Thus $\mu^{*}(v)\geq 4-6\times\frac{1}{3}-4\times\frac{1}{3}-\frac{1}{4}-3\times\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R6(b) and R7, and at most $\frac{1}{4}$ to its $3$-, $4(1)$-, $4(2)$- and $5(4^{+})$-neighbours by R3-R5 and R8-R9, $\frac{1}{18}$ to each of its $6(5)$-neighbours by R10.
@@
(6.3). Let $m_{3}(v)=7$. Notice first that if $v$ has at most five $5^{-}$-neighbours, then we have $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and at most $\frac{1}{3}$ to each of its $5^{-}$-neighbours by R4-R9. Therefore we may assume that $v$ has at least six $5^{-}$-neighbours, which clearly infers that $v$ has at least one support vertex. Thus we have $d_{2}(v)\geq 23$ by Lemma 2.38. Observe that $v$ has at most six $4$-neighbours as $d_{2}(v)\geq 23$. In particular, by Lemma 2.29(c), a $4(3^{+})$-vertex has no $4$- and $5(4^{+})$-neighbours. So, we deduce that $v$ has at most four $4(3^{+})$-neighbours. If $v$ has four such neighbours, then $v$ is not adjacent to any $4(2^{-})$- or $5(4^{+})$-vertices by Lemma 2.29(c). Thus $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-4\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R6(b) and R7. We may further assume that $v$ has at most three $4(3^{+})$-neighbours.
Suppose first that $v$ has no $4(3^{+})$-neighbour. Note that $v$ has at most six neighbours consisting of $4(2^{-})$- or 5(5)555(5)5 ( 5 )-vertices as $d_{2}(v)\geq 23$. If $v$ has six such neighbours, then $v$ cannot be adjacent to any $5(4)$-vertex as $d_{2}(v)\geq 23$, and so $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-6\times\frac{1}{4}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, at most $\frac{1}{4}$ to each of its $4(1)$-, $4(2)$- and 5(5)555(5)5 ( 5 )-neighbours by R4-R5 and R9(b). If $v$ has at most five neighbours consisting of $4(2^{-})$- or 5(5)555(5)5 ( 5 )-vertices, then $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-5\times\frac{1}{4}-3\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, at most $\frac{1}{4}$ to each of its $4(1)$-, $4(2)$- and 5(5)555(5)5 ( 5 )-neighbours by R4-R5 and R9(b), $\frac{1}{9}$ to each of its $5(4)$-neighbours by R8.
Next we suppose that $v$ has exactly one $4(3^{+})$-neighbour. By Lemma 2.29(c), $v$ has at least one neighbour different from $4$- and $5(4^{+})$-vertices. On the other hand, a $4(t)$-vertex with $1\leq t\leq 2$ has at most one neighbour consisting of $4$- or 5(5)555(5)5 ( 5 )-vertices by Lemma 2.29(a),(b). It then follows from Lemma 2.29(c) and Lemma 2.34(b) that $v$ has at most four neighbours consisting of $4(1)$-, $4(2)$- or 5(5)555(5)5 ( 5 )-vertices. Thus $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-\frac{1}{3}-4\times\frac{1}{4}-2\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R6(b) and R7, at most $\frac{1}{4}$ to each of its $4(1)$-, $4(2)$- and 5(5)555(5)5 ( 5 )-neighbours by R4-R5 and R9(b), $\frac{1}{9}$ to each of its $5(4)$-neighbours by R8.
Suppose that $v$ has exactly two $4(3^{+})$-neighbours. By Lemma 2.29(c), a $4(3^{+})$-vertex has neither $4$-neighbour nor $5(4^{+})$-neighbour. This means that $v$ has at least two neighbours different from $4$- and $5(4^{+})$-vertices. Thus $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-2\times\frac{1}{3}-4\times\frac{1}{4}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R6(b) and R7, and at most $\frac{1}{4}$ to each of its $4(1)$-, $4(2)$- and $5(4^{+})$-neighbours by R4-R5 and R8-R9.
Finally suppose that $v$ has exactly three $4(3^{+})$-neighbours. It follows from Lemma 2.29(c) that $v$ has at least three neighbours different from $4$- and $5(4^{+})$-vertices. Thus $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-3\times\frac{1}{3}-2\times\frac{1}{4}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R6(b) and R7, at most $\frac{1}{4}$ to each of its $4(1)$-, $4(2)$- and $5(4^{+})$-neighbours by R4-R5 and R8-R9.

### v3

@@ 2.3. The case $\Delta=8$
(5.3). Let $m_{3}(v)=6$. Notice first that if $v$ has at most four neighbours forming $4$- or $5$-vertices, then we have $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-4\times\frac{2}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and at most $\frac{2}{9}$ to each of its $4(1)$-, $4(2)$-, $4(3)$-, and $5(4^{+})$-neighbours by R7. Therefore we may further assume that $v$ has at least five neighbours forming $4$- or $5$-vertices. By Lemma 2.31, we conclude that a $4(1^{+})$-vertex has at most one $4$-neighbour. In addition, a $5(5)$-vertex has at most one neighbour consisting of $4$- or $5(5)$-vertices by Lemma 2.37(b). This implies that $v$ has at most four $5(5)$-neighbours; in particular, $v$ has at most five $4(1^{+})$-neighbours.
@@
Suppose first that $v$ has no $5(5)$-neighbours. If $v$ has at most four $4(1^{+})$-neighbours, then $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-4\times\frac{1}{6}-3\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{6}$ to each of its $4(1^{+})$-neighbours by R7(a), and $\frac{1}{9}$ to each of its $5(4)$-neighbours by R7(b). If $v$ has five $4(1^{+})$-neighbours, then $v$ cannot have any $5(4)$-neighbours by Lemma 2.40(b). So, $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-5\times\frac{1}{6}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{1}{6}$ to each of its $4(1^{+})$-neighbours by R7(a).
@@
Next we suppose that $v$ has exactly one $5(5)$-neighbour. By Lemma 2.40(c), $v$ has at most four $4$-neighbours. If $v$ has at most two $4$-neighbours, then $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-\frac{2}{9}-2\times\frac{1}{6}-4\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{2}{9}$ to its $5(5)$-neighbour by R7(c), $\frac{1}{6}$ to each of its $4(1^{+})$-neighbours by R7(a), and $\frac{1}{9}$ to each of its $5(4)$-neighbours by R7(b). If $v$ has at least three $4$-neighbours, then $v$ is adjacent to at most one $5(4)$-vertices by Lemma 2.40(c), and so $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-\frac{2}{9}-4\times\frac{1}{6}-\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{2}{9}$ to its $5(5)$-neighbour by R7(c), $\frac{1}{6}$ to each of its $4(1^{+})$-neighbours by R7(a), and $\frac{1}{9}$ to each of its $5(4)$-neighbours by R7(b).
@@
Suppose now that $v$ has exactly two $5(5)$-neighbours. Then, by Lemma 2.40(d), $v$ has at most three neighbours consisting of $4$- or $5(4)$-vertices, and so $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-2\times\frac{2}{9}-3\times\frac{1}{6}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{2}{9}$ to each of its $5(5)$-neighbours by R7(c), and at most $\frac{1}{6}$ to each of its $4(1^{+})$ and $5(4)$-neighbours by R7(a),(b).
@@
Suppose that $v$ has exactly three $5(5)$-neighbours. It follows that $v$ has at most one $4$-neighbour by Lemma 2.40(e). By the same reason, $v$ has at most two neighbours consisting of $4$- or $5(4)$-vertices. Thus, we have $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-3\times\frac{2}{9}-\frac{1}{6}-\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{2}{9}$ to its $5(5)$-neighbour by R7(c), $\frac{1}{6}$ to each of its $4(1^{+})$-neighbours by R7(a), and $\frac{1}{9}$ to each of its $5(4)$-neighbours by R7(b).
@@
Finally, suppose that $v$ has four $5(5)$-neighbours. Then $v$ has neither $4$- nor $5(4)$-neighbours by Lemma 2.37(a),(b), and so $\mu^{*}(v)\geq 3-6\times\frac{1}{3}-4\times\frac{2}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{2}{9}$ to each of its $5(5)$-neighbours by R7(c).
@@
(5.4). Let $m_{3}(v)=7$. By Lemma 2.40(f), $v$ has at most five $5(4^{+})$-neighbours, and so $\mu^{*}(v)\geq 3-7\times\frac{1}{3}-5\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-face by R1, $\frac{1}{9}$ to each of its $5(4^{+})$-neighbour by R8.
@@
(6). Let $k=8$. The initial charge of $v$ is $\mu(v)=d(v)-4=4$. Notice first that if $m_{3}(v)\leq 4$, then we have $\mu^{*}(v)\geq 4-4\times\frac{1}{3}-8\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and at most $\frac{1}{3}$ to each of its neighbours by R3, R5, R9. Therefore, we may assume that $5\leq m_{3}(v)\leq 8$.
@@
(6.1). Let $m_{3}(v)=5$. By Lemma 2.31(c), a $4(3^{+})$-vertex cannot be adjacent to any $4$-vertex. It follows that $v$ has at most four $4(3^{+})$-neighbours. Then $\mu^{*}(v)\geq 4-5\times\frac{1}{3}-4\times\frac{1}{3}-4\times\frac{1}{4}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R9(c), and at most $\frac{1}{4}$ to each of its other neighbours by R3, R5, R9.
@@
(6.2). Let $m_{3}(v)=6$. By Lemma 2.31(c), $v$ has at most four $4(3^{+})$-neighbours. Suppose first that $v$ has no $4(3^{+})$-neighbour. Then $\mu^{*}(v)\geq 4-6\times\frac{1}{3}-8\times\frac{1}{4}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and at most $\frac{1}{4}$ to each of its neighbours by R3, R5, R9.
@@
Next suppose that $v$ has exactly one $4(3^{+})$-neighbour. It follows from Lemmas 2.30(a) and 2.31(c) that $v$ has at most six neighbours consisting of $3$-, $4(2^{-})$- or $5(4^{+})$-vertices. Thus $\mu^{*}(v)\geq 4-6\times\frac{1}{3}-\frac{1}{3}-6\times\frac{1}{4}-\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to its $4(3^{+})$-neighbour by R9(c), at most $\frac{1}{4}$ to each of its $3$-, $4(1)$-, $4(2)$- and $5(4^{+})$-neighbours by R3, R9, and $\frac{1}{18}$ to each of its $6(5)$-neighbours by R5.
@@
Suppose now that $v$ has exactly two $4(3^{+})$-neighbours. In such a case, $v$ has at most five neighbours consisting of $3$-, $4(2^{-})$- or $5(4^{+})$-vertices by Lemmas 2.30(a) and 2.31(c). Thus $\mu^{*}(v)\geq 4-6\times\frac{1}{3}-2\times\frac{1}{3}-5\times\frac{1}{4}-\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R9(c), at most $\frac{1}{4}$ to each of its $3$-, $4(1)$-, $4(2)$- and $5(4^{+})$-neighbours by R3, R9, and $\frac{1}{18}$ to each of its $6(5)$-neighbours by R5.
@@
Suppose that $v$ has exactly three $4(3^{+})$-neighbours. In such a case, $v$ can have at most three neighbours consisting of $3$-, $4(2^{-})$- or $5(4^{+})$-vertices by Lemmas 2.30(a) and 2.31(c). Thus $\mu^{*}(v)\geq 4-6\times\frac{1}{3}-3\times\frac{1}{3}-3\times\frac{1}{4}-2\times\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R9(c), at most $\frac{1}{4}$ to each of its $3$-, $4(1)$-, $4(2)$- and $5(4^{+})$-neighbours by R3, R9, and $\frac{1}{18}$ to each of its $6(5)$-neighbours by R5.
@@
Finally suppose that $v$ has exactly four $4(3^{+})$-neighbours. In such a case, $v$ can have at most one neighbour consisting of $3$-, $4(2^{-})$- or $5(4^{+})$-vertices by Lemmas 2.30(a) and 2.31(c). Thus $\mu^{*}(v)\geq 4-6\times\frac{1}{3}-4\times\frac{1}{3}-\frac{1}{4}-3\times\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R9(c), and at most $\frac{1}{4}$ to its $3$-, $4(1)$-, $4(2)$- and $5(4^{+})$-neighbours by R3, R9, and $\frac{1}{18}$ to each of its $6(5)$-neighbours by R5.
(6.3). Let $m_{3}(v)=7$. Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[7]$. Clearly, $v$ has no $3$-neighbour by Lemma 2.30.
@@
Let us first claim that $v$ has a support neighbour. By contradiction, assume that $v$ does not have. Then, $v_{2},v_{3},\ldots,v_{7}$ are $5^{+}$-vertices. Recall that a $5(5)$-vertex has at most one $5(5)$-neighbour by Lemma 2.37(b). Thus we infer that $v$ has at most four $5(5)$-neighbour. If one of $v_{1},v_{8}$ is a $5^{+}$-vertex, then $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-\frac{1}{3}-4\times\frac{2}{9}-3\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, at most $\frac{1}{3}$ to each of its $4$-neighbours by R9(a)-(c), $\frac{2}{9}$ to each of its $5(5)$-neighbours by R9(e), $\frac{1}{9}$ to each of its $5(4)$-neighbours by R9(d). Thus, we assume that $v_{1}$ and $v_{8}$ are $4$-vertices. By Lemma 2.31, this infer that $v$ has at most three $5(5)$-neighbours. It then follows that $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-2\times\frac{1}{3}-3\times\frac{1}{9}-3\times\frac{2}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, at most $\frac{1}{3}$ to each of its $4$-neighbours by R9(a)-(c), $\frac{1}{9}$ to each of its $5(4)$-neighbours by R9(d), $\frac{2}{9}$ to each of its $5(5)$-neighbours by R9(e). Consequently, we deduce that $v$ has a support neighbour, i.e., one of $v_{2},v_{3},\ldots,v_{7}$ is a $4^{-}$-vertex.